# Query-Based Adaptive Aggregation for Multi-Dataset Joint Training Toward Universal Visual Place Recognition

### v1

@@ Abstract
Deep learning methods for Visual Place Recognition (VPR) have advanced significantly, largely driven by large-scale datasets. However, most existing approaches are trained on a single dataset, which can introduce dataset-specific inductive biases and limit model generalization. While multi-dataset joint training offers a promising solution for developing universal VPR models, divergences among training datasets can saturate limited information capacity in feature aggregation layers, leading to suboptimal performance. To address these challenges, we propose Query-based Adaptive Aggregation (QAA), a novel feature aggregation technique that leverages learned queries as reference codebooks to effectively enhance information capacity without significant computational or parameter complexity. We show that computing the Cross-query Similarity (CS) between query-level image features and reference codebooks provides a simple yet effective way to generate robust descriptors. Our results demonstrate that QAA outperforms state-of-the-art models, achieving balanced generalization across diverse datasets while maintaining peak performance comparable to dataset-specific models. Ablation studies further explore QAA’s mechanisms and scalability. Visualizations reveal that the learned queries exhibit diverse attention patterns across datasets. Code will be publicly released.
Introduction
Visual Place Recognition (VPR) (Schubert et al. 2024; Zaffar et al. 2021; Berton et al. 2022) is a fundamental computer vision task that involves retrieving the top-K most similar images from a database of geo-referenced or pose-annotated images given a query image. Learning-based VPR methods (Arandjelovic et al. 2016; Ge et al. 2020; Berton et al. 2023; Berton, Masone, and Caputo 2022; Izquierdo and Civera 2024a, b; Ali-bey, Chaib-draa, and Giguère 2022, 2023; Lu et al. 2024; Ali-bey, Chaib-draa, and Giguère 2024; Tzachor et al. 2025; Xiao, Zhu, and Loianno 2025; Leyva-Vallina, Strisciuglio, and Petkov 2023; Trivigno et al. 2023) are widely applied in computer vision and robotics, particularly in tasks such as large-scale coarse-to-fine camera localization (Sarlin et al. 2019), loop closure in SLAM systems (Masone and Caputo 2021; Campos et al. 2021), and absolute localization in GPS-denied environments (Couturier and Akhloufi 2021; Xiao et al. 2023). However, outdoor VPR still faces challenges, including significant variations due to domain shifts (e.g., day-night variation), changes in viewpoint (e.g., front-view vs. multi-view), occlusion by moving objects, and the absence of prominent landmarks. To address these challenges, researchers have developed several outdoor large-scale VPR datasets (Warburg et al. 2020; Berton, Masone, and Caputo 2022; Ali-bey, Chaib-draa, and Giguère 2022) to train robust models, each capturing different environmental conditions and scene characteristics. These datasets introduce specific biases influenced by factors such as camera viewpoint, domain, sampling density, and geographic diversity, outlined in Table 1.
@@ Introduction
Joint training across multiple VPR datasets (Izquierdo and Civera 2024a; Berton and Masone 2025; Liu et al. 2025) has emerged as a promising approach to achieving more robust and universally applicable outdoor VPR models. One pioneering example, SALAD Clique Mining (CM) (Izquierdo and Civera 2024a), introduced a clique-based grouping method to cluster densely sampled images from the MSLS dataset (Warburg et al. 2020), facilitating joint training with the sparser GSV-Cities dataset (Ali-bey, Chaib-draa, and Giguère 2022). Our cross-dataset evaluations similarly indicate that models trained exclusively on single datasets become biased toward specific dataset characteristics, resulting in limited generalization capability. Conversely, joint training across multiple diverse datasets consistently yields superior performance in a broad range of benchmarks. Nonetheless, our results also show that the baseline aggregation protocol proposed in (Izquierdo and Civera 2024b) occasionally underperforms relative to dataset-specific models, likely due to significant training dataset divergence and limited information capacity in feature aggregation layers.
To address these challenges, we introduce Query-based Adaptive Aggregation (QAA) (Fig. 1), a novel feature aggregation method aimed at enhancing multi-dataset joint training performance. QAA employs learned queries as reference codebooks to efficiently expand memory within aggregation layers, thereby increasing information capacity and maintaining strong performance even with low-dimensional descriptors. This approach strengthens cross-domain generalization while preserving peak performance comparable to dataset-specific models. By utilizing the cross-query similarity matrix between query-level image features and the independent reference codebook, QAA effectively models robust geographic descriptors that capture the relative spatial relationships between images. The adaptive learned queries enable the generation of diverse feature representations across datasets while introducing minimal computational and parameter overhead. The main contributions are as follows:
@@
  • We introduce Cross-query Similarity (CS), a simple yet effective aggregation paradigm that constructs similarity matrices between image features and reference codebooks to generate robust geographic descriptors. We analyze its information capacity through coding rate (Ma et al. 2007; Yu et al. 2020), providing insights into its superior performance.
@@
Learning-based VPR methods can be broadly categorized into one-stage (Arandjelovic et al. 2016; Ge et al. 2020; Berton et al. 2023; Berton, Masone, and Caputo 2022; Izquierdo and Civera 2024a, b; Ali-bey, Chaib-draa, and Giguère 2022, 2023; Leyva-Vallina, Strisciuglio, and Petkov 2023) and two-stage approaches (Zhu et al. 2023; Wang et al. 2022; Lu et al. 2024; Tzachor et al. 2025; Trivigno et al. 2023). One-stage methods typically employ a CNN (LeCun, Bengio, and Hinton 2015) or ViT (Dosovitskiy et al. 2021) model as the feature extractor, generating local feature maps, followed by a feature aggregation module to convert the 2D feature map into a 1D global descriptor. The query image descriptor is then compared against a database to identify the top-K similar images and retrieve their geo-referenced information. Two-stage methods, in contrast, re-rank these top-k candidates by further utilizing the 2D feature maps, achieving enhanced localization performances with higher computational costs. In VPR, the output dimensionality of descriptors plays a crucial role, with a linear impact on both memory usage and matching cost.
@@
Score-based feature aggregation methods (Arandjelovic et al. 2016; Izquierdo and Civera 2024b), also known as cluster-based aggregation, compute global descriptors by weighting and summing patch-level image features based on predicted scores. For instance, NetVLAD (Arandjelovic et al. 2016) uses softmax scores to weight the distances between image features and cluster centroids. Similarly, SALAD (Izquierdo and Civera 2024b) weights patch features by scores normalized with the Sinkhorn optimal transport algorithm (Sinkhorn and Knopp 1967; Cuturi 2013). In contrast, our QAA method eliminates explicit score prediction. Instead, it generates robust descriptors by computing a cross-query similarity matrix between learned query-level image features and an independent reference codebook, achieving comparable or reduced dimensionality.
@@
Figure 2: The training framework (top) and Query-Based Adaptive Aggregation (QAA) architecture (bottom)
@@
Recent VPR methods leverage foundational models like DINOv2 (Oquab et al. 2024) for better generalization. For example, AnyLoc (Keetha et al. 2023) leverages general-purpose representations from DINOv2, SelaVPR adds lightweight adapters to DINOv2, and EffoVPR uses DINOv2’s self-attention features for reranking. The most relevant work, BoQ (Ali-bey, Chaib-draa, and Giguère 2024), utilizes learned queries to project image features from a DINOv2 backbone. Unlike BoQ’s concatenation approach, which combines all learned query outputs and reduces dimensionality with linear layers, our QAA method computes cross-query similarity with independent reference codebooks. This approach prevents the output dimension from increasing with the number of queries, enabling scalable query use with a fixed output size.
Methodology
Framework Overview
@@
We present the training framework (Fig. 2) for multi-dataset training. We denote $\mathcal{D}=\{D_{1},D_{2},\cdots,D_{n}\}$ as training datasets, where $n$ is the number of datasets. From each dataset, we sample $k$ places, with each place consisting of $m$ images. The concatenated batch from all datasets is represented as =$\mathcal{B}=\{B_{1},B_{2},\cdots,B_{n}\}$, where $\mathcal{B}$ contains $n\times k\times m$ images in total as the input image batch.
@@
We use DINOv2-B/14 (Oquab et al. 2024) as the backbone to balance performance and latency. The input has dimensions $N\times C\times H\times W$, where $N=n\times k\times m$ represents the number of images per batch, and $C$, $H$, and $W$ are the image channels, height, and width. The output includes a feature map $X$ of size $N\times P\times C_{o}$, with $P=\frac{H}{14}\times\frac{W}{14}$ and $C_{o}$ as the output feature map channels. We then employ the Query-based Adaptive Aggregation (QAA) method for feature aggregation. The final outputs $\mathcal{E}$ are descriptors of size $N\times C_{d}$, where $C_{d}$ represents the output descriptor dimension. This output can be decomposed as =$\mathcal{E}=\{E_{1},E_{2},\cdots,E_{n}\}$ for each dataset. The overall framework design allows the flexibility to incorporate additional datasets for joint training.
@@
We employ Multi-Similarity (MS) Loss (Wang et al. 2019; Ali-bey, Chaib-draa, and Giguère 2022) to train the final blocks of the DINOv2 model along with the QAA module. Images from the same place are grouped and treated as positive examples, while images from different places (groups) serve as negative samples.
Query-based Adaptive Aggregation (QAA)
@@
Query-Level Image Features. To generate query-level image features $\hat{P}$, QAA employs two modules: 1) a Feature Self-Attention (Feature Self-Attn) mechanism, utilizing a Multi-Head Attention (MHA) module (Vaswani et al. 2017), and 2) a Feature Prediction module, which integrates an MHA module with a projection layer to reduce the channel dimension to $C_{f}$, where $C_{f}$ represents the number of channels for $\hat{P}$. This process is defined as:
@@
[…]
where $\hat{P}$ denotes the query-level image features with dimensions $N_{q}\times C_{f}$ (omitting the batch dimension for simplicity) and $\hat{Q}^{f}$ denotes the self-attention-refined version of $Q^{f}$. The patch-level image feature $X$ serves as the key and value in the feature prediction module. The inclusion of the self-attention module, inspired by (Ali-bey, Chaib-draa, and Giguère 2024), minimizes the need for substantial modifications to $Q^{f}$, thereby enhancing training stability and accelerating convergence. $\hat{Q}^{f}$ can be cached after training.
@@
where $S$ represents the pairwise similarity matrix between $\hat{F}$ and $\hat{P}$ along the query dimension of $N_{q}$, justifying the term Cross-query Similarity. Notably, this mechanism is similar to the similarity computation between keys and queries in the attention mechanism (Vaswani et al. 2017). However, a key distinction is that while attention computes similarity along the channel dimension, QAA performs similarity computation along the query dimension. Equation 4 fundamentally computes the cross-correlation matrix between query-level image features and codebooks, effectively capturing second-order statistics along the query dimension.
The final descriptor $\mathcal{E}$ is obtained by applying intra-L2 normalization on $S$ along the $C_{r}$ dimension (Arandjelovic et al. 2016), followed by a global L2 normalization of the entire vector. The output descriptor dimension $C_{d}$ is given by $C_{r}\times C_{f}$. During inference, the aggregation process involves only the computations from Equations 2 and 4, along with the final normalization. This ensures that QAA maintains minimal computational and parameter complexity.
Our intuition behind the CS paradigm is to preserve the information capacity of $\hat{P}$. Unlike score-based paradigms that compress the output space into the range $[0,1]$, the CS paradigm avoids such projection, retaining more information for interaction with the reference codebook. To quantify the information retained in $\hat{P}$, we leverage the coding rate concept from information theory (Ma et al. 2007; Yu et al. 2020). The coding rate is computed as:
$R(\hat{P}^{\top},\epsilon)=\frac{1}{2}\log\det\left(I+\frac{C_{f}}{N_{q}\epsilon^{2}}\hat{P}^{\top}\hat{P}\right)$ (5)
@@
CS aggregation paradigm represents a significant advancement in retrieval-based VPR, as it demonstrates—for the first time—that informative geographic descriptors can be formulated directly from the similarity matrix between query-level image features and an independent reference codebook. This eliminates the need for explicit score prediction techniques (Arandjelovic et al. 2016; Izquierdo and Civera 2024b) or implicit linear projection. Therefore, CS aggregation enhances the interpretability of descriptor generation in retrieval-based VPR.
Table 1: Overview of large-scale VPR datasets for joint training.  Sampling density refers to the spacing between sampled locations, with each location potentially containing multiple images.
Datasets GSV-Cities (Ali-bey, Chaib-draa, and Giguère 2022) MSLS (Warburg et al. 2020) SF-XL (Berton, Masone, and Caputo 2022)
@@
Table 2: Recall@1 Comparison of state-of-the-art VPR methods with our results on large-scale multi-view VPR datasets. The best results are highlighted in bold and the rest of the top-3 results are underlined.
@@
NetVLAD (Arandjelovic et al. 2016) VGG-16 4096 16.3 77.7 85.9 85.0 - 40.0 76.9 69.8
SRFS (Ge et al. 2020) VGG-16 4096 29.7 72.3 90.4 89.1 - 50.3 83.8 80.3
Conv-AP (Ali-bey, Chaib-draa, and Giguère 2022) ResNet-50 4096 33.9 87.5 92.4 90.5 80.1 47.5 74.4 76.2
MixVPR (Ali-bey, Chaib-draa, and Giguère 2023) ResNet-50 4096 40.2 89.4 94.2 91.5 85.2 71.1 88.5 85.1
CosPlace (Berton, Masone, and Caputo 2022) ResNet-50 2048 47.7 90.0 92.3 90.9 75.3 76.4 88.8 87.3
EigenPlace (Berton et al. 2023) ResNet-50 2048 48.9 90.7 94.1 92.5 82.4 84.1 90.8 93.0
BoQ (Ali-bey, Chaib-draa, and Giguère 2024) DINOv2-B 12288 63.0 92.2 96.6 93.7 92.5 91.8 95.2 98.1
SALAD CM (Izquierdo and Civera 2024b, a) DINOv2-B 8448 58.1 91.9 95.2 92.6 89.1 85.6 94.6 96.8
@@
Table 3: Recall@1 Comparison of state-of-the-art VPR methods with our results on front-view VPR datasets. The best results are highlighted in bold and the rest of the top-3 results are underlined.
@@
NetVLAD (Arandjelovic et al. 2016) VGG-16 4096 58.9 - - 13.1 8.0 66.4 51.5 54.4 35.4
SRFS (Ge et al. 2020) VGG-16 4096 70.0 - - 16.0 28.6 81.1 69.7 76.0 54.8
Conv-AP (Ali-bey, Chaib-draa, and Giguère 2022) ResNet-50 4096 83.4 - 38.2 62.9 43.4 91.9 82.8 91.0 80.4
MixVPR (Ali-bey, Chaib-draa, and Giguère 2023) ResNet-50 4096 88.0 64.0 58.4 76.2 64.4 96.2 91.5 96.8 84.8
CosPlace (Berton, Masone, and Caputo 2022) ResNet-50 2048 87.4 67.5 54.4 71.9 50.7 92.2 87.0 92.0 78.5
EigenPlace (Berton et al. 2023) ResNet-50 2048 89.2 67.4 54.2 71.2 58.9 93.1 90.0 93.1 86.4
BoQ (Ali-bey, Chaib-draa, and Giguère 2024) DINOv2-B 12288 93.8 79.0 81.3 90.6 97.7 98.5 98.8 99.4 97.5
SALAD CM (Izquierdo and Civera 2024b, a) DINOv2-B 8448 94.2 82.7 90.7 95.2 95.6 98.5 98.4 99.2 98.1

### v4

@@ Abstract
Deep learning methods for Visual Place Recognition (VPR) have advanced significantly, largely driven by large-scale datasets. However, most existing approaches are trained on a single dataset, which can introduce dataset-specific inductive biases and limit model generalization. While multi-dataset joint training offers a promising solution for developing universal VPR models, divergences among training datasets can saturate the limited information capacity in feature aggregation layers, leading to suboptimal performance. To address these challenges, we propose Query-based Adaptive Aggregation (QAA), a novel feature aggregation technique that leverages learned queries as reference codebooks to effectively enhance information capacity without significant computational or parameter complexity. We show that computing the Cross-query Similarity (CS) between query-level image features and reference codebooks provides a simple yet effective way to generate robust descriptors. Our results demonstrate that QAA outperforms state-of-the-art models, achieving balanced generalization across diverse datasets while maintaining peak performance comparable to dataset-specific models. Ablation studies further explore QAA’s mechanisms and scalability. Visualizations reveal that the learned queries exhibit diverse attention patterns across datasets. Project page: xjh19971.github.io/QAA.
I Introduction
Visual Place Recognition (VPR) [29, 44, 7] is a fundamental robotic perception task that involves retrieving the top-K most similar images from a database of geo-referenced or pose-annotated images given a query image. Learning-based VPR methods [4, 16, 9, 5, 17, 18, 1, 2, 24, 3, 35, 41, 21, 34] are widely applied in robotics and computer vision, particularly in tasks such as large-scale coarse-to-fine camera localization [28], loop closure in SLAM systems [26, 10], and absolute localization in GPS-denied environments [12, 40]. However, outdoor VPR still faces challenges, including significant variations due to domain shifts (e.g., day-night variation), changes in viewpoint (e.g., front-view vs. multi-view), occlusion by moving objects, and the absence of prominent landmarks. To address these challenges, researchers have developed several outdoor large-scale VPR datasets [39, 5, 1] to train robust models, each capturing different environmental conditions and scene characteristics. These datasets introduce specific biases influenced by factors such as camera viewpoint, domain, sampling density, and geographic diversity (Table I).
@@ I Introduction
Joint training across multiple VPR datasets [17, 6, 22] has emerged as a promising approach to achieving more robust and universally applicable outdoor VPR models. One pioneering example, SALAD Clique Mining (CM) [17], introduced a clique-based grouping method to cluster densely sampled images from the MSLS dataset [39], facilitating joint training with the sparser GSV-Cities dataset [1]. Our cross-dataset evaluations similarly indicate that models trained exclusively on single datasets become biased toward specific dataset characteristics, resulting in limited generalization capability. Conversely, joint training across multiple diverse datasets consistently yields superior performance in a broad range of benchmarks. Nonetheless, our results also show that the baseline aggregation protocol proposed in [18] occasionally underperforms relative to dataset-specific models, likely due to significant training dataset divergence and limited information capacity in feature aggregation layers.
To address these challenges, we introduce Query-based Adaptive Aggregation (QAA) (Fig. 1), a novel feature aggregation method aimed at enhancing multi-dataset joint training performance. QAA employs learned queries as reference codebooks to efficiently expand memory within aggregation layers, thereby increasing information capacity and maintaining strong performance even with low-dimensional descriptors. This approach strengthens cross-domain generalization while preserving peak performance comparable to dataset-specific models. By utilizing the cross-query similarity matrix between query-level image features and the independent reference codebook, QAA effectively models robust geographic descriptors that capture the relative spatial relationships between images. The adaptive learned queries enable the generation of diverse feature representations across datasets while introducing minimal computational and parameter overhead. The main contributions are:
@@
  • We introduce Cross-query Similarity (CS), a simple yet effective aggregation paradigm that constructs similarity matrices between image features and reference codebooks to generate robust geographic descriptors. We analyze its information capacity through coding rate [25, 43], providing insights into its performance.
@@ II Related Works
Learning-based VPR methods can be broadly categorized into one-stage [4, 16, 9, 5, 17, 18, 1, 2, 21] and two-stage approaches [45, 37, 24, 35, 34]. One-stage methods typically employ a CNN [20] or ViT [15] model as the feature extractor, generating local feature maps, followed by a feature aggregation module to convert the 2D feature map into a 1D global descriptor. The query image descriptor is then compared against a database to identify the top-K similar images and retrieve their geo-referenced information. Two-stage methods, in contrast, re-rank these top-k candidates by further utilizing the 2D feature maps, achieving enhanced localization performances with higher computational costs. In VPR, the output dimensionality of descriptors plays a crucial role, with a linear impact on both memory usage and matching cost.
@@
Score-based feature aggregation methods [4, 18], also known as cluster-based aggregation, compute global descriptors by weighting and summing patch-level image features based on predicted scores. For instance, NetVLAD [4] uses softmax scores to weight the distances between image features and cluster centroids. Similarly, SALAD [18] weights patch features by scores normalized with the Sinkhorn optimal transport algorithm [30, 14]. In contrast, our QAA method eliminates explicit score prediction. Instead, it generates robust descriptors by computing a cross-query similarity matrix between learned query-level image features and an independent reference codebook, achieving comparable or reduced dimensionality.
@@
Figure 2: The training framework (top) and Query-Based Adaptive Aggregation (QAA) architecture (bottom).
@@
Recent VPR methods leverage foundational models like DINOv2 [27] for better generalization. For example, AnyLoc [19] leverages general-purpose representations from DINOv2, SelaVPR adds lightweight adapters to DINOv2, and EffoVPR uses DINOv2’s self-attention features for reranking. The most relevant work, BoQ [3], utilizes learned queries to project image features from a DINOv2 backbone. Unlike BoQ’s concatenation approach, which combines all learned query outputs and reduces dimensionality with linear layers, our QAA method computes cross-query similarity with independent reference codebooks. This approach prevents the output dimension from increasing with number of queries, enabling scalable query use with a fixed output size.
III Methodology
III-A Framework Overview
@@
We present the training framework (Fig. 2) for multi-dataset training. We denote $\mathcal{D}=\{D_{1},D_{2},\cdots,D_{n}\}$ as training datasets, where $n$ is the number of datasets. From each dataset, we sample $k$ places, with each place consisting of $m$ images. The concatenated batch from all datasets is represented as $\mathcal{B}=\{B_{1},B_{2},\cdots,B_{n}\}$, where $\mathcal{B}$ contains $n\times k\times m$ images in total as the input image batch.
@@
We use DINOv2-B/14 [27] as the backbone to balance performance and latency. The input has dimensions $N\times C\times H\times W$, where $N=n\times k\times m$ represents the number of images per batch, and $C$, $H$, and $W$ are the image channels, height, and width. The output includes a feature map $X$ of size $N\times P\times C_{o}$, with $P=\frac{H}{14}\times\frac{W}{14}$ and $C_{o}$ as the output feature map channels. We then employ the Query-based Adaptive Aggregation (QAA) method for feature aggregation. The final outputs $\mathcal{E}$ are descriptors of size $N\times C_{d}$, where $C_{d}$ represents the output descriptor dimension. This output can be decomposed as $\mathcal{E}=\{E_{1},E_{2},\cdots,E_{n}\}$ for each dataset. The overall framework design allows the flexibility to incorporate additional datasets for joint training.
@@
We employ Multi-Similarity (MS) Loss [38, 1] to train the final blocks of the DINOv2 model along with the QAA module. Images from the same place are grouped and treated as positive examples, while images from different places (groups) serve as negative samples.
III-B Query-based Adaptive Aggregation (QAA)
@@
Query-Level Image Features. To generate query-level image features $\hat{P}$, QAA employs two modules: 1) a Feature Self-Attention (Feature Self-Attn) mechanism, utilizing a Multi-Head Attention (MHA) module [36], and 2) a Feature Prediction module, which integrates an MHA module with a projection layer to reduce the channel dimension to $C_{f}$, where $C_{f}$ represents the number of channels for $\hat{P}$. This process is defined as:
@@
$\hat{P}$ $=\text{Feature-Prediction}(\hat{Q}^{f},X,X),$ (2)
where $\hat{P}$ denotes the query-level image features with dimensions $N_{q}\times C_{f}$ (omitting the batch dimension) and $\hat{Q}^{f}$ denotes the self-attention-refined version of $Q^{f}$. The patch-level image feature $X$ serves as the key and value in the feature prediction module. The inclusion of the self-attention module, inspired by [3], minimizes the need for substantial modifications to $Q^{f}$, thereby enhancing training stability and accelerating convergence. $\hat{Q}^{f}$ can be cached after training.
@@
where $S$ represents the pairwise similarity matrix between $\hat{F}$ and $\hat{P}$ along the query dimension of $N_{q}$, justifying the term Cross-query Similarity. Notably, this mechanism is similar to the similarity computation between keys and queries in the attention mechanism [36]. However, a key distinction is that while attention computes similarity along the channel dimension, QAA performs similarity computation along the query dimension. Equation 4 fundamentally computes the cross-correlation matrix between query-level image features and codebooks, effectively capturing second-order statistics along the query dimension.
The final descriptor $\mathcal{E}$ is obtained by applying intra-L2 normalization on $S$ along the $C_{r}$ dimension [4], followed by a global L2 normalization of the entire vector. The output descriptor dimension $C_{d}$ is given by $C_{r}\times C_{f}$. During inference, the aggregation process involves only the computations from Equations 2 and 4, along with the final normalization. This ensures that QAA maintains minimal computational and parameter complexity.
Our intuition behind the CS paradigm is to preserve the information capacity of $\hat{P}$. Unlike score-based paradigms that compress the output space into the range $[0,1]$, the CS paradigm avoids such projection, retaining more information for interaction with the reference codebook. To quantify the information retained in $\hat{P}$, we leverage the coding rate concept from information theory [25, 43]. The coding rate is computed as:
[…]
@@
CS aggregation paradigm represents a significant advancement in retrieval-based VPR, as it demonstrates—for the first time—that informative geographic descriptors can be formulated directly from the similarity matrix between query-level image features and an independent reference codebook. This eliminates the need for explicit score prediction techniques [4, 18] or implicit linear projection. Therefore, CS aggregation enhances the interpretability of descriptor generation in retrieval-based VPR.
TABLE I: Overview of large-scale VPR datasets for joint training. Sampling density refers to the spacing between sampled locations, with each location potentially containing multiple images.
Datasets GSV-Cities [1] MSLS [39] SF-XL [5]
@@
TABLE II: Recall@1 Comparison of State-of-the-Art VPR Methods with Our Results on Multi-view VPR Datasets. The best results are highlighted in bold and the rest of the top-3 results are underlined.
@@
NetVLAD [4] VGG-16 4096 16.3 77.7 85.9 85.0 - 40.0 76.9 69.8
SRFS [16] VGG-16 4096 29.7 72.3 90.4 89.1 - 50.3 83.8 80.3
Conv-AP [1] ResNet-50 4096 33.9 87.5 92.4 90.5 80.1 47.5 74.4 76.2
MixVPR [2] ResNet-50 4096 40.2 89.4 94.2 91.5 85.2 71.1 88.5 85.1
CosPlace [5] ResNet-50 2048 47.7 90.0 92.3 90.9 75.3 76.4 88.8 87.3
EigenPlace [9] ResNet-50 2048 48.9 90.7 94.1 92.5 82.4 84.1 90.8 93.0
BoQ [3] DINOv2-B 12288 63.0 92.2 96.6 93.7 92.5 91.8 95.2 98.1
SALAD CM [18, 17] DINOv2-B 8448 58.1 91.9 95.2 92.6 89.1 85.6 94.6 96.8
@@
TABLE III: Recall@1 Comparison of State-of-the-Art VPR Methods with Our Results on Front-view VPR Datasets. The best results are highlighted in bold and the rest of the top-3 results are underlined.
@@
NetVLAD [4] VGG-16 4096 58.9 - - 13.1 8.0 66.4 51.5 54.4 35.4
SRFS [16] VGG-16 4096 70.0 - - 16.0 28.6 81.1 69.7 76.0 54.8
Conv-AP [1] ResNet-50 4096 83.4 - 38.2 62.9 43.4 91.9 82.8 91.0 80.4
MixVPR [2] ResNet-50 4096 88.0 64.0 58.4 76.2 64.4 96.2 91.5 96.8 84.8
CosPlace [5] ResNet-50 2048 87.4 67.5 54.4 71.9 50.7 92.2 87.0 92.0 78.5
EigenPlace [9] ResNet-50 2048 89.2 67.4 54.2 71.2 58.9 93.1 90.0 93.1 86.4
BoQ [3] DINOv2-B 12288 93.8 79.0 81.3 90.6 97.7 98.5 98.8 99.4 97.5
SALAD CM [18, 17] DINOv2-B 8448 94.2 82.7 90.7 95.2 95.6 98.5 98.4 99.2 98.1